# Logarithmic Smoothing for Adaptive PAC-Bayesian Off-Policy Learning

### v1

@@ 3 Warm-up on Pessimistic Logarithmic Smoothing
3.1 Batch learning algorithms with theoretical guarantees
Pessimistic Logarithmic Smoothing recently appeared as a fruitful theoretical approach to obtain both concrete learning algorithms and generalization guarantees guarantees for contextual bandits. Those coupled benefits naturally stems from the use of PAC-Bayesian theory alquier2024user which provide empirical generalization bounds whom the optimization yield original learning algorithms with theoretical guarantees. In sakhi2024logarithmic , pessimistic PAC-Bayes bounds have been developed for contextual bandits, providing a novel learning algorithm with optimality guarantee. We first recall the definition of Logarithmic smoothing as well as its associated regularized IPS estimator. For any $\lambda>0,k\geq 0$,
$h_{\mathrm{LS}}^{\lambda\mathrm{-Lin}}(p,q,c):=-\frac{p}{\lambda}\log\left(1-\frac{\lambda c}{q}\right)\quad\text{and}\quad\hat{R}_{0:k}^{\lambda\mathrm{-LS}}:=\hat{R}_{0:k}^{h_{\mathrm{LS}}^{\lambda\mathrm{-Lin}}}.$
Using these notions, sakhi2024logarithmic obtained results for static off-policy learning i.e. data are gathered in a single batch $S_{0}$ of size $n_{0}$. We recall them below.
Proposition 3.1 (Propositions 10 and 11 of sakhi2024logarithmic, ).
Given a prior $P\in\mathcal{P}(\Theta),\delta\in(0,1]$ and $\lambda>0$, the following holds with probability at least $1-\delta$ over $S_{0}$:
$\forall Q\in\mathcal{P}(\Theta),\quad R\left(\pi_{Q}\right)\leq\hat{R}_{0}^{\lambda-\mathrm{LS}}\left(\pi_{Q}\right)+\frac{\operatorname{KL}(Q\|P)+\log\frac{1}{\delta}}{\lambda n_{0}}.$
Furthermore, for any $\mathcal{C}(\Theta)\subset\mathcal{P}(\Theta)$, denoting by $Q_{n_{0}}\in\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}\hat{R}_{0}^{\lambda-\mathrm{LS}}\left(\pi_{Q}\right)+\frac{\operatorname{KL}(Q\|P)}{\lambda n_{0}}$ and $Q^{\star}\in\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}R(\pi_{Q})$, we have:
00 0 $\leq R(\pi_{Q_{n_{0}}})-R(\pi_{Q^{\star}})\leq\lambda\mathcal{S}(\pi_{Q^{\star}})+\frac{2\operatorname{KL}(Q^{\star}\|P)+\log\frac{2}{\delta}}{\lambda n_{0}},$
@@ 3.1 Batch learning algorithms with theoretical guarantees
Those results provide two distinct hindsights: (i), the learning procedure yielding $Q_{n_{0}}$ (and exploited in practice in sakhi2024logarithmic ) minimises a generalisation bound and (ii), once $\mathcal{S}(\pi_{Q^{\star}})$ finite, taking $\lambda=1/\sqrt{n_{0}}$ gives a convergence rate of $1/\sqrt{n_{0}}$ to the optimal value of the risk. However, Proposition 3.1 and its associated learning algorithm suffers from theoretical and practical limitations.
@@
    A batch procedure cannot handle a progressive arrival of data $(S_{k})_{k\geq 0}$. In particular, Proposition 3.1 does not allow to re-train our policy w.r.t all past data when a new data batch arrives.
@@
    The pseudo-variance term of Proposition 3.1 does not go to zero, even in the oracle case where $\pi_{0}=\pi_{Q^{\star}}$, whatever the optimal policy $\pi_{Q^{\star}}$ is. This exhibits a fundamental limitation of this batch procedure as the slow convergence rate of $1/\sqrt{n_{0}}$ is inevitable.
We provide below alternatives for those two fundamental issues. To answer 1. we show in Sec. 3.2 that leveraging online PAC-Bayes and supermartingale techniques haddouche2022online ; haddouche2023pac it is possible to obtain an adaptive version of Proposition 3.1, allowing both the arrival of new data on-the-fly and re-using former ones. Then, re-using online PAC-Bayes techniques, we answer 2. in Sec. 4 by providing a novel adaptive learning algorithm with accelerated convergence rate.
3.2 Towards a first adaptive pessimistic learning algorithm
In this part, we assume, for pedagogical concerns, that all datasets $S_{k}$ have the same size $n_{k}=m$. Still exploiting Logarithmic Smoothing, we provide an adaptive counterpart of Proposition 3.1 by involving online PAC-Bayes learning techniques haddouche2022online ; haddouche2023pac . From a technical perpective, we allow the progressive arrival of data batches $(S_{k})_{k\geq 0}$ by controlling the expansion of a supermartingale through Ville’s inequality doob1939jean instead of an exponential moment via Markov’s inequality. Our result is stated as follows:
@@ 3.2 Towards a first adaptive pessimistic learning algorithm
theoremextensionsakhi For any adaptive algorithm Alg, any $m>0$ and data independent prior $P$. For any $\lambda>0,\delta\in(0,1]$, with probability at least $1-\delta$ over $S$, we have, for all $k\geq 0,$ denoting by $\pi_{k+1}$ the output of Alg at time $k$, for all $Q\in\mathcal{P}(\Theta)$:
@@
The proof lies in Appendix D. Note that taking $k=0$ and $m=n_{0}$ retrieves exactly the batch case of Proposition 3.1. On the contrary taking $k=n$ and $m=1$ provides a purely online procedure. Our theorem provides an interpolation between batch and online scenarios, allowing sub-batches of data to arrive at each time step and optimize on-the-fly.
Notice that the optimization problem $\operatorname{argmin}_{Q\in\mathcal{P}(\Theta)}\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{Q})+\frac{\operatorname{KL}(Q,P)}{\lambda(k+1)m}$ admits a closed form solution: the Gibbs posterior $\mathrm{Q}_{\lambda}$ such that $dQ_{\lambda}(\theta)\propto\exp\left(-\lambda(k+1)m\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(d_{\theta})\right)$ (catoni2007pac, , Theorem 1.2.6). However, Gibbs posteriors often requires costly, advanced Monte Carlo methods mcbook ; smc_book to be computed. A common practice to approximate the Gibbs posterior is to restrict the optimization problem on a subclass $\mathcal{C}(\Theta)\subset\mathcal{P}(\Theta)$ on which the optimisation problem is tractable, a widely used class in PAC-Bayes learning is the class of Gaussian distributions dziugaite2017computing ; perez2021tighter ; haddouche2022online .
Finally, a key remark is that, given a fixed batch size $m>0$ and parameter $\lambda>0$, Sec. 3.2 holds for a countable number $k$ of sub-batches simultaneously. A fundamental difference with Proposition 3.1 is that we allow the empirical risk $\hat{R}_{0:k}^{\lambda-\mathrm{LS}}$ to evolve with $k$, while the batch approach only tackles $S_{0}$. This is crucial as $\hat{R}_{0:k}^{\lambda-\mathrm{LS}}$ that evolve adaptively to the intermediate policies $(\pi_{j})_{j=0\cdots k}$ while $\hat{R}_{0}^{\lambda-\mathrm{LS}}$ only involves the initial policy $\pi_{0}$, which does improve during the learning process. Thus, optimising Sec. 3.2 at each time step yields an adaptive pessimistic learning algorithm given in Algorithm 1.
1 Input: Behavior policy $\pi_{0}$, a policy class $\mathcal{C}(\Theta)$, $\lambda>0$.
@@ 5       Solve the optimization problem:
$\hat{Q}_{k+1}=\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}\left\{\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{\mathrm{Q}})+\frac{\operatorname{KL}(Q,P)}{\lambda km}\right\}$ (1)
@@
Algorithm 1 Adaptive Policy Learning via Logarithmic Smoothing
@@
Note that at time $k$, we instantiate $\hat{R}_{0:k}^{\lambda-\mathrm{LS}}$ with $(\pi_{j})_{j=0\cdots k}$ the outputs of Algorithm 1 up to time $k-1$. Similarly to the batch case, Algorithm 1 also enjoys a convergence guarantee. {restatable}propositionslowconvergence For any fixed subbatch size $m>0$, any data independent prior $P$, any $\lambda>0,\delta\in(0,1]$, with probability at least $1-\delta$, for any $k\geq 0$, denoting by $\pi_{k+1}$ the output of Algorithm 1 at time $k$ and $\pi_{\star}=\pi_{Q_{\star}}$ with $Q_{\star}\in\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}R(\pi_{Q})$:
@@
$R\left(\pi_{k+1}\right)-R\left(\pi_{\star}\right)$ $\leq\frac{\lambda}{k+1}\sum_{j=0}^{k}\mathcal{S}_{j}\left(\pi_{\star}\right)+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)m}$
@@
where $\mathcal{S}_{j}(\pi):=\mathbb{E}_{(x,a,c)\sim P(\pi_{j})}\left[\frac{\pi(a\mid x)c^{2}}{\pi_{j}^{2}(a\mid x)}\right]$. Algorithm 1 is proven in Appendix D, and taking $\lambda=\frac{1}{\sqrt{(k+1)m}}$ yields an overall convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{km}}\right)$, matching the rate of Proposition 3.1 in the batch case (i.e $k=0,m=n_{0}$) and yielding an original rate of $\mathcal{O}(1/\sqrt{n})$ for an online setting ($k=n,m=1$). Algorithm 1 also works for any intermediate regime between batch and online settings, showing the strength of the adaptive approach. Looking at Algorithm 1, we observe that the upper bound depends on the average $\frac{1}{k+1}\sum_{j=0}^{k}\mathcal{S}_{j}(\pi_{\star})$, which is expected to be smaller than $\mathcal{S}_{0}(\pi_{\star})$, as intermediate policies $\pi_{j}$ should be progressively closer to $\pi_{\star}$. However, this algorithm still suffers from the same limitations than the static case here as the pseudo-variance terms $\mathcal{S}_{j}$ do not decrease to zero, even in the perfect situation where for all $j$, $\pi_{j}=\pi_{\star}$.
@@
In this part, we showed that online PAC-Bayes techniques successfully extend offline algorithms to adaptive ones, improving dependencies while enjoying the same convergence rates. A key question is then: is it possible to find an adaptive algorithm with an accelerated convergence rate?
4 Adaptive Logarithmic Smoothing with accelerated convergence.
While Sec. 3.2 provide a natural adaptive extension of sakhi2024logarithmic , Algorithm 1 still suffers (Algorithm 1) from a slow convergence rate of $\mathcal{O}(1/\sqrt{km})$ due to the nature of the pseudo-variance term. Is it then possible to avoid such a term? Yes, by adjusting $h^{\lambda-\mathrm{Lin}}$ introduced in Sec. 3: we define $h^{\text{adj}}_{\lambda}(p,q,c):=-\frac{p}{\lambda}\log\left(1-\frac{\lambda c}{q(1+\lambda c)}\right)$ and denote it as the adjusted linear LS regularizer. We then define the adjusted risk for any adaptive algorithm Alg and $\lambda\in(0,1),k\geq 0$:
$\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi):=R_{0,k}^{h^{\text{adj}}_{\lambda}}(\pi)=-\frac{1}{N_{k}}\sum_{j=1}^{k}\sum_{i\in I_{j}}\frac{\pi(a_{i}|x_{i})}{\lambda}\log\left(1-\frac{\lambda c_{i}}{\pi_{j}(a_{i}|x_{i})(1+\lambda c_{i})}\right),$ (2)
where again, $\pi_{k+1}$ is the output of Alg at time $k$. Using similar tools than Sec. 3.2, we are able to reach the following theoretical guarantee for our novel regularizer $h^{\text{adj}}_{\lambda}$
Theorem 4.1.
For any adaptive algorithm Alg, any sequence $(n_{k})_{k\geq 0}$ and data independent prior $P$. For any $\boldsymbol{\lambda\in(0,1)},\delta\in(0,1]$, with probability at least $1-\delta$ over $S$, we have, for all $k\geq 0$, denoting by $\pi_{k+1}$ the output of Alg at time $k$, for all $Q\in\mathcal{P}(\Theta)$:
$R\left(\pi_{Q}\right)-\sum_{j=0}^{k}\frac{n_{j}}{N_{k}}R\left(\pi_{j}\right)$ $\leq\hat{R}^{\lambda-\mathrm{adj}}_{0:k}(\pi_{Q})+\hat{C}_{0,k}(\lambda)+\frac{\operatorname{KL}(Q,P)+\log\left(\frac{1}{\delta}\right)}{\lambda N_{k}},$
Thm. 4.1 is obtained in Appendix E, using the same tools than Sec. 3.2 and thus enjoys the same general remarks. Note that this result holds for varying batch sizes $(n_{k})_{k\geq 0}$, allowing to encompass real-life situations where there are discrepancies in the arrival of new data. In particular, it is the theoretical backbone behind the main algorithm of this work: an adaptive and adjusted learning algorithm for contextual bandits, stated in Algorithm 2.
1 Input: Behavior policy $\pi_{0}$, a policy class $\mathcal{C}(\Theta)$, $\lambda\in(0,1)$.

### v2

@@ 3 WARM-UP ON PESSIMISTIC LOGARITHMIC SMOOTHING
3.1 Background: Offline algorithms with theoretical guarantees
PAC-Bayesian Logarithmic Smoothing recently appeared as a fruitful theoretical approach to obtain both concrete learning algorithms and generalization guarantees for contextual bandits. In Sakhi et al. (2024), PAC-Bayesian bounds have been developed for off-policy contextual bandits, and optimizing them provide novel learning algorithms with optimality guarantees. We first recall the definition of the logarithmic smoothing estimator (LS) as well as its associated regularizer. For any $\lambda>0$ we define LS for $k\geq 0$,
$h_{\mathrm{LS}}^{\lambda\mathrm{-Lin}}(p,q,c):=-\frac{p}{\lambda}\log\left(1-\frac{\lambda c}{q}\right)\,$,\,
$\hat{R}_{0:k}^{\lambda\mathrm{-LS}}:=\hat{R}_{0:k}^{h_{\mathrm{LS}}^{\lambda\mathrm{-Lin}}}.$
Using these notions, Sakhi et al. (2024) obtained results for static off-policy learning i.e. data are gathered in a single batch $S_{0}$ of size $n_{0}$. We recall them below.
Proposition 3.1 (Propositions 10 and 11 of Sakhi et al., 2024).
Given a prior $P\in\mathcal{P}(\Theta),\delta\in(0,1]$ and $\lambda>0$, the following holds with probability at least $1-\delta$ over $S_{0},\forall Q\in\mathcal{P}(\Theta),$ :
$R\left(\pi_{Q}\right)\leq\hat{R}_{0}^{\lambda-\mathrm{LS}}\left(\pi_{Q}\right)+\frac{\operatorname{KL}(Q\|P)+\log\frac{1}{\delta}}{\lambda n_{0}}.$
Furthermore, for any $\mathcal{C}(\Theta)\subset\mathcal{P}(\Theta)$, denoting by $\hat{Q}\in\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}\hat{R}_{0}^{\lambda-\mathrm{LS}}\left(\pi_{Q}\right)+\frac{\operatorname{KL}(Q\|P)}{\lambda n_{0}}$ and $Q^{\star}\in\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}R(\pi_{Q})$, we have:
$R(\pi_{\hat{Q}})-R(\pi_{Q^{\star}})\leq\lambda\mathcal{S}(\pi_{Q^{\star}})+\frac{2\operatorname{KL}(Q^{\star}\|P)+\log\frac{2}{\delta}}{\lambda n_{0}},$
@@ 3.1 Background: Offline algorithms with theoretical guarantees
Those results provide two distinct insights: (i), the learning procedure yielding $\hat{Q}$ (and exploited in practice in Sakhi et al. (2024)) minimizes a generalization bound and (ii), once $\mathcal{S}(\pi_{Q^{\star}})$ finite, taking $\lambda=1/\sqrt{n_{0}}$ gives a convergence rate of $1/\sqrt{n_{0}}$ to the optimal value of the risk. However, Proposition˜3.1 and its associated learning algorithm suffer from theoretical and practical limitations.
@@
    A batch procedure cannot handle a progressive arrival of data $(S_{k})_{k\geq 0}$. In particular, Proposition˜3.1 does not allow to retrain our policy w.r.t all past data when a new data batch arrives.
@@
    The pseudo-variance term of Proposition˜3.1 does not go to zero, even in the oracle case where $\pi_{0}=\pi_{Q^{\star}}$, whatever the optimal policy $\pi_{Q^{\star}}$ is. This issue is not an artifact of the proof technique, but from the performance objective $R(\pi)$, and consequently its estimator. This exhibits a fundamental limitation of this procedure as the slow convergence rate of $1/\sqrt{n_{0}}$ is inevitable.
We answer those two fundamental issues below. To answer 1, we show in Sec.˜3.2 that leveraging online PAC-Bayes techniques allows obtaining a sequential version of Proposition˜3.1, permitting both the arrival of new data on-the-fly and naturally re-using former ones. Then, we answer 2, in Sec.˜4 by adjusting the performance objective and its respective estimator, providing a novel sequential learning algorithm with accelerated convergence.
3.2 Towards a first sequential algorithm
In this part, we assume, for pedagogical concerns, that all datasets $S_{k}$ have the same size $n_{k}=m$. Still exploiting Logarithmic Smoothing, we provide a sequential counterpart of Proposition˜3.1 by involving online PAC-Bayes learning techniques (Haddouche and Guedj, 2022, 2023). Our result is stated as follows:
@@ 3.2 Towards a first sequential algorithm
theoremextensionsakhi For any sequential algorithm Alg, any $m>0$ and data independent prior $P$. For any $\lambda>0,\delta\in(0,1]$, with probability at least $1-\delta$ over $S$, we have, for all $k\geq 0,$ denoting by $\pi_{k+1}$ the output of Alg at time $k$, for all $Q\in\mathcal{P}(\Theta)$:
@@
The proof lies in Appendix˜B. Note that taking $k=0$ and $m=n_{0}$ retrieves exactly the batch case of Proposition˜3.1. On the contrary taking $k=n$ and $m=1$ provides a purely online procedure. Our theorem provides an interpolation between batch and online scenarios, allowing sub-batches of data to arrive at each time step and optimize on-the-fly.
1Input: Policy $\pi_{0}$, policy class $\mathcal{C}(\Theta)$, $\lambda>0$.
@@ 5   Solve the optimization problem:
$\hskip-5.69054pt\hat{Q}_{k+1}=\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}\left\{\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{\mathrm{Q}})+\frac{\operatorname{KL}(Q,P)}{\lambda(k+1)m}\right\}$
@@
Algorithm 1 Sequential Policy Learning via Logarithmic Smoothing
While Sec.˜3.2 has a similar shape to Proposition˜3.1, the involved toolbox differs fundamentally. Indeed, while batch PAC-Bayes only requires controlling an exponential moment via Markov’s inequality, Online PAC-Bayes requires a careful design and control of a supermartingale through Ville’s inequality (Doob, 1939). This core difference allows, beyond Sec.˜3.2, to reach a more general result, stated in Appendix˜C, and allowing any arbitrary evolution of the context distribution. In particular, we can avoid the i.i.d. assumption on contexts made in Sec.˜2. This better reflects real-world applications, where user behavior and responses naturally evolve over time. Although analyzing algorithm convergence in this non-stationary setting is of interest, it falls outside the scope of the present work.
Notice that the optimization problem $\operatorname{argmin}_{Q\in\mathcal{P}(\Theta)}\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(\pi_{Q})+\frac{\operatorname{KL}(Q,P)}{\lambda(k+1)m}$ admits a closed form solution: the Gibbs posterior $\mathrm{Q}_{\lambda}$ such that $dQ_{\lambda}(\theta)\propto\exp\left(-\lambda(k+1)m\hat{R}_{0:k}^{\lambda-\mathrm{LS}}(d_{\theta})\right)$ (Catoni, 2007, Theorem 1.2.6). However, Gibbs posteriors often require costly, advanced Monte Carlo methods (Chopin and Papaspiliopoulos, 2020) to be computed. A common practice is to restrict the optimization problem on a subclass $\mathcal{C}(\Theta)\subset\mathcal{P}(\Theta)$ on which the optimization problem is tractable, e.g. the class of Gaussian distributions (Dziugaite and Roy, 2017; Perez-Ortiz et al., 2021). Restricting the family of distributions is not limiting: as long as the family of parametrized predictors $\{f_{\theta}\mid\theta\in\Theta\}$ is expressive enough, the induced policy family still contains the optimal policy (Sakhi et al., 2023a).
Finally, a key remark is that, given a fixed batch size $m>0$ and parameter $\lambda>0$, Sec.˜3.2 holds for a countable number $k$ of sub-batches simultaneously. A fundamental difference with Proposition˜3.1 is that we allow the empirical risk $\hat{R}_{0:k}^{\lambda-\mathrm{LS}}$ to evolve with $k$, while the batch approach only tackles a single $S_{0}$. This allows the optimization of the right-hand-side of Sec.˜3.2 at each time step and yields a sequential learning algorithm given in Algorithm˜1.
@@
In Algorithm˜1, at time $k$, we instantiate $\hat{R}_{0:k}^{\lambda-\mathrm{LS}}$ with $(\pi_{j})_{j=0\cdots k}$ the outputs of Algorithm˜1 up to time k1k-1. Similarly to the batch case, Algorithm˜1 also enjoys a convergence guarantee to the optimal policy. {restatable}propositionslowconvergence For any fixed subbatch size $m>0$, any data independent prior $P$, any $\lambda>0,\delta\in(0,1]$, with probability at least $1-\delta$, for any $k\geq 0$, denoting by $\pi_{k+1}$ the output of Algorithm˜1 at time $k$ and $\pi_{\star}=\pi_{Q_{\star}}$ with $Q_{\star}\in\operatorname{argmin}_{Q\in\mathcal{C}(\Theta)}R(\pi_{Q})$:
@@
$R\left(\pi_{k+1}\right)-R\left(\pi_{\star}\right)\\ \leq\frac{\lambda}{k+1}\sum_{j=0}^{k}\mathcal{S}_{j}\left(\pi_{\star}\right)+2\frac{\operatorname{KL}(Q^{\star},P)+\log\left(\frac{2}{\delta}\right)}{\lambda(k+1)m}$
where $\mathcal{S}_{j}(\pi):=\mathbb{E}_{(x,a,c)\sim P(\pi_{j})}\left[\frac{\pi(a\mid x)c^{2}}{\pi_{j}^{2}(a\mid x)}\right]$. Algorithm˜1 is proven in Appendix˜B. To obtain explicit convergence rates, we make the assumption that our policy class covers the optimal actions, i.e.:
Assumption 3.2 (Coverage of the optimal actions).
For any policy $\pi$ in our policy class $\mathcal{C}(\Theta)$, we assume that: $\inf_{x}\pi(a^{\star}(x)|x)\geq C^{\star}$ for some constant $C^{\star}>0$.
@@
Assumption˜3.2 only requires coverage of the optimal actions, which is strictly weaker than the uniform coverage condition assumed in prior work (Zenati et al., 2023). With this assumption, we ensure that $S_{j}(\pi_{\star})$ is finite and uniformly upper bounded by $1/C^{\star}$ for all $j$. Taking $\lambda=\frac{1}{\sqrt{(k+1)m}}$ yields an overall convergence rate of $\mathcal{O}\left(\frac{1}{\sqrt{(k+1)m}}\right)$, matching the rate of Proposition˜3.1 in the batch case (i.e $k=0,m=n_{0}$) and yielding an original rate of $\mathcal{O}(1/\sqrt{n})$ for an online setting ($k=n,m=1$). Algorithm˜1 also works for any intermediate regime between batch and online settings, showing the strength of the sequential approach. Looking at Algorithm˜1, we observe that the upper bound depends on the average $\frac{1}{k+1}\sum_{j=0}^{k}\mathcal{S}_{j}(\pi_{\star})$. Intuitively, this average is expected to be smaller than $\mathcal{S}_{0}(\pi_{\star})$, as intermediate policies $\pi_{j}$ should be progressively closer to $\pi_{\star}$. However, this algorithm still suffers from the same limitations to the static case here as the pseudo-variance terms $\mathcal{S}_{j}$ do not decrease to zero, even in the perfect situation where for all $j$, $\pi_{j}=\pi_{\star}$.
@@
In this part, we showed that online PAC-Bayes techniques successfully extend offline algorithms to the sequential setting, revealing the dependence on the data-collection process while enjoying the same convergence rates. A key question is then: is it possible to find a sequential algorithm with a faster convergence rate?
4 SEQUENTIAL LOGARITHMIC SMOOTHING WITH ACCELERATED CONVERGENCE
1Input: Policy $\pi_{0}$, a policy class $\mathcal{C}(\Theta)$, $\lambda\in(0,1)$.